OpenStudy (anonymous):

Find the equation of the line that passes through the points (0, 8) and (-5, 4).

OpenStudy (anonymous):

the form of equation of every line is : y=a*x+b. in order to find a and b u have to substitute x and y value of the points given. for example in your case: 8=a*0+b from that u can find b=8. and from the other point passing the line we have: 4=a*(-5)+b , b now is found so 4=a*(-5)+8 from this eq u can finf the value of a